# Towards domain-invariant Self-Supervised Learning with Batch Styles Standardization

### 21 Dec 2023 v5

@@ 8 $r\sim U(r_{min},r_{max})$
%**** iclr2024_conference_main.tex Line 700 ****the minimum and maximum possible
@@
of the target amplitudes.
%**** iclr2024_conference_main.tex Line 725 ****# Compute center coordinates of amplitudes
@@
n_views (int): Number of augmented views
@@
%**** iclr2024_conference_main.tex Line 775 ****amp = self.substitute_low_freq(
@@ B.1 SimCLR
$\mathcal{L}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}=\displaystyle\dfrac{-1}{V-1}\sum_{s^{\prime}\neq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}\log\left(\dfrac{e^{{\bm{z}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}\cdot{\bm{z}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c}s^{\prime}}/T}}{\displaystyle\sum_{(c^{\prime\prime},s^{\prime\prime})\neq({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s})}e^{{\bm{z}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}\cdot{\bm{z}}_{c^{\prime\prime}s^{\prime\prime}}/T}}\right)$ (6)
@@ B.2 SWaV
To compute cluster assignments also referred to as codes and denoted ${\bm{q}}_{n}^{(v)}$, SWaV relies on the Sinkhorn-Klopp (SK) algorithm (Cuturi, 2013). SK is performed on all views representations trying to assign representations to the most similar centroids but also uniformly among clusters. Finally, based on the swapped prediction concept, SWaV minimizes the following per-sample loss:
@@
In practice, SWaV employs a multi-crop strategy, generating $2$ global views (large crops) and $V$ local views (small crops) for each image. Cluster assignments are then computed only from the $2$ global views while probabilities are derived from all the $V+2$ views. In this setting, SWaV minimizes the following loss:
@@ B.3 MSN
In practice and more formally, MSN generates for each image ${\bm{X}}_{n}$, $M$ masked views $\{{\bm{X}}_{n,1},\ldots,{\bm{X}}_{n,M}\}$ and a single unmasked view ${\bm{X}}_{n}^{+}$. Masked views are processed by a student encoder and the unmasked view by a teacher encoder whose weights are updated via an exponential moving average of the student encoder’s weights. Masked and unmasked views’ embeddings denoted $\{{\bm{z}}_{n,1},\ldots,{\bm{z}}_{n,M}\},{\bm{z}}_{n}^{+}$ are then compared to a set of centroids/prototypes ${\bm{C}}\in\mathbb{R}^{K\times D}$ and the resulting similarities are converted into probability distributions $\{{\bm{p}}_{n,1},\ldots,{\bm{p}}_{n,M}\},{\bm{p}}_{n}^{+}$:
@@ C.1 Pretraining
On PACS and DomainNet, as part of the geometric augmentations, we use random crop resizing, horizontal flips, small rotations, cutout(DeVries & Taylor, 2017) while color augmentations are applied in batch-wise manner using color jitter, random equalize, random posterize, random solarize and random grayscale. On Camelyon17 WILDS, we use random crop resizing, flips, rotations, cutout and batch-wise color jitter. All other hyperparameters for SimCLR, SWaV, MSN are respectively specified on Tables 5,  6,  7.
@@
For all datasets (PACS, DomainNet, and Camelyon17 WILDS), we use the Adam optimization method (Kingma & Ba, 2014) with an initial learning rate of $10^{-4}$, a learning rate scheduler with cosine decay, and weight decay of $10^{-4}$. The networks are trained respectively for $5$K, $1$K, and $15$K steps with batch sizes of $128$, $64$, and $64$. When performing linear probing, we follow the same normalization scheme as (He et al., 2022) by adding a batch normalization layer (Ioffe & Szegedy, 2015) without affine parameters before the linear classifier.
@@ D.1 Transfer learning in DG/UDG
The usage of Transfer Learning in DG/UDG is common but we think it is misguided. Often the pretraining dataset, such as Imagenet, can include one or more of the target domains, e.g., photo for PACS or real for DomainNet. When evaluating on these domains, it is not possible to know if performances result from the generalization ability of the DG/UDG methods or from the transfer learning. For new DG/UDG methods, it is hard not to follow the common practice because transfer learning unfairly boosts the results of previous works, and state-of-the-art performances are often seen as a prerequisite for paper acceptance. We tried to limit the usage of transfer learning in our experiments and only used it for the DomainNet dataset. The bias introduced by Imagenet transfer learning can be seen in our results and especially for the domains that are close to Imagenet: on PACS (Table 1 main paper), without transfer learning, SimCLR with BSS is consistently the best method for all settings except for the photo domain, where methods using Imagenet transfer learning report better performance. However, on DomainNet (Table 2 main paper), when using Imagenet transfer learning similar to the other UDG methods, SimCLR with BSS achieves better performances on the real domain for 2 out of 3 proportions of labeled data ($1\%$, $5\%$). In the $10\%$ labeled data setting, where SimCLR with BSS is outperformed, it is worth noting that DiMAE is the only method to perform full fine-tuning which is probably the reason behind the gap in performances.
@@ D.3 Impact of $r$ on BSS generated images
To illustrate the effect of the hyperparameter $r$ on the generated images by BSS, we apply BSS on a single batch fixing the chosen style image and varying $r$. The resulting images are reported in Figure 5. We can observe that as $r$ increases, textures/styles with higher frequencies are transferred to the resulting images.

### 19 Jan 2024 v6

@@ 8 $r\sim U(r_{min},r_{max})$
the minimum and maximum possible
@@
%**** iclr2024_conference_supp.tex Line 175 ****of the target amplitudes.
# Compute center coordinates of amplitudes
@@
%**** iclr2024_conference_supp.tex Line 200 ****n_views (int): Number of augmented views
@@
amp = self.substitute_low_freq(
@@ B.1 SimCLR
$\mathcal{L}_{{\color[rgb]{0,0,0}c}{\color[rgb]{0,0,0}s}}=\displaystyle\dfrac{-1}{V-1}\sum_{s^{\prime}\neq{\color[rgb]{0,0,0}s}}\log\left(\dfrac{e^{{\bm{z}}_{{\color[rgb]{0,0,0}c}{\color[rgb]{0,0,0}s}}\cdot{\bm{z}}_{{\color[rgb]{0,0,0}c}s^{\prime}}/T}}{\displaystyle\sum_{(c^{\prime\prime},s^{\prime\prime})\neq({\color[rgb]{0,0,0}c},{\color[rgb]{0,0,0}s})}e^{{\bm{z}}_{{\color[rgb]{0,0,0}c}{\color[rgb]{0,0,0}s}}\cdot{\bm{z}}_{c^{\prime\prime}s^{\prime\prime}}/T}}\right)$ (1)
@@ B.2 SWaV
To compute cluster assignments also referred to as codes and denoted ${\bm{q}}_{n}^{(v)}$, SWaV relies on the Sinkhorn-Klopp (SK) algorithm (cuturi2013sinkhorn). SK is performed on all views representations trying to assign representations to the most similar centroids but also uniformly among clusters. Finally, based on the swapped prediction concept, SWaV minimizes the following per-sample loss:
@@
In practice, SWaV employs a multi-crop strategy, generating $2$ global views (large crops) and $V$ local views (small crops) for each image. Cluster assignments are then computed only from the $2$ global views while probabilities derived from all the $V+2$ views. In this setting, SWaV minimizes the following loss:
@@ B.3 MSN
In practice and more formally, MSN generates for each image ${\bm{X}}_{n}$, $M$ masked views $\{{\bm{X}}_{n,1},\ldots,{\bm{X}}_{n,M}\}$ and a single unmasked view ${\bm{X}}_{n}^{+}$. Masked views are processed by a student encoder and the unmasked view by teacher encoder whose weights are updated via exponential moving average of the student encoder’s weights. Masked and unmasked views’ embeddings denoted $\{{\bm{z}}_{n,1},\ldots,{\bm{z}}_{n,M}\},{\bm{z}}_{n}^{+}$ are then compared to a set of centroids/prototypes ${\bm{C}}\in\mathbb{R}^{K\times D}$ and the resulting similarities are converted into probability distributions $\{{\bm{p}}_{n,1},\ldots,{\bm{p}}_{n,M}\},{\bm{p}}_{n}^{+}$:
@@ C.1 Pretraining
On PACS and DomainNet, as part of the geometric augmentations, we use random crop resizing, horizontal flips, small rotations, cutout(devries2017improved) while color augmentations are applied in batch-wise manner using color jitter, random equalize, random posterize, random solarize and random grayscale. On Camelyon17 WILDS, we use random crop resizing, flips, rotations, cutout and batch-wise color jitter. All other hyperparameters for SimCLR, SWaV, MSN are respectively specified on Tables 1,  2,  3.
@@
For all datasets (PACS, DomainNet, and Camelyon17 WILDS), we use the Adam optimization method (kingma2014adam) with an initial learning rate of $10^{-4}$, a learning rate scheduler with cosine decay, and weight decay of $10^{-4}$. The networks are trained respectively for $5$K, $1$K, and $15$K steps with batch sizes of $128$, $64$, and $64$. When performing linear probing, we follow the same normalization scheme as (he2022masked) by adding a batch normalization layer (ioffe2015batch) without affine parameters before the linear classifier.
@@ D.1 Transfer learning in DG/UDG
The usage of Transfer Learning in DG/UDG is common but we think it is misguided. Often the pretraining dataset, such as Imagenet, can include one or more of the target domains, e.g., photo for PACS or real for DomainNet. When evaluating on these domains, it is not possible to know if performances result from the generalization ability of the DG/UDG methods or from the transfer learning. For new DG/UDG methods, it is hard not to follow the common practice because transfer learning unfairly boosts the results of previous works, and state-of-the-art performances are often seen as a prerequisite for paper acceptance. We tried to limit the usage of transfer learning in our experiments and only used it for the DomainNet dataset. The bias introduced by Imagenet transfer learning can be seen in our results and especially for the domains that are close to Imagenet: on PACS (Table 1 main paper), without transfer learning, SimCLR with BSS is consistently the best method for all settings except for the photo domain, where methods using Imagenet transfer learning report better performance. However, on DomainNet (Table 2 main paper), when using Imagenet transfer learning similar to the other UDG methods, SimCLR with BSS achieves better performances on the real domain for 2 out of 3 proportions of labeled data ($1\%$, $5\%$). In the $10\%$ labeled data setting, where SimCLR with BSS is outperformed, it is worth noting that DiMAE is the only method to perform full fine-tuning which is probably the reason behind the gap of performances.
@@ D.3 Impact of $r$ on BSS generated images
To illustrate the effect of the hyperparamter $r$ on the generated images by BSS, we apply BSS on a single batch fixing the chosen style image and varying $r$. The resulting images are reported in Figure 2. We can observe that as $r$ increases, textures/styles with higher frequencies are transferred to the resulting images.